If secA – tanA = p
Then find the value of secA, tanA.

Grade:11

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:Hello student, please find answer to your question
secA - tanA = p….....(1)
\frac{1}{cosA} - \frac{sinA}{cosA} = p
\frac{(secA - tanA)(secA+tanA)}{(secA + tanA)} = p
\frac{(sec^{2}A - tan^{2}A)}{(secA + tanA)} = p
\frac{1}{(secA + tanA)} = p
secA + tanA = \frac{1}{p}…......(2)
(1) + (2)
2secA =p+ \frac{1}{p}
secA =\frac{1}{2}(p+ \frac{1}{p})
(1) – (2)
tanA =\frac{1}{2}(p- \frac{1}{p})